Wednesday 08/28/12

9:50 A-7:00P, 8:30P – 11:50P

  • pick up drive from HCBI
  • STORM4 lost focus O/N.  Reset focus, continue imaging.
  • Trying to get good matches for pnrPr dynamics in simulation, not just end states.
  • Dyes not switching well in 15 min etched K27me3.  Especially A647.  Also has obvious glowing chunks.
  • Etching K27 Pc large slide from ~4/20.  20 min etch, still solidly embedded. Etching another 20 min.
  • fixed bug in lsm_tif_2_mytif.
  • 6 rounds of 20 min etches, K27 Pc etching complete.  Manual check confirms nice A647 switching.
  • Tiling K27 Pc large slide for O/N STORM imaging.
  • Solved issue with 647 not switching –> 405 was up to 10 mW, completely largely bleached 647 channel prior to imaging during the 750 run.  Set down to 1.2 mW.  Recalibrating Pc K27.  Should re-image the Pc-GFP… 🙁
  • interestingly almost same proportional increases in dot intensity with the 10x weaker 405.
  • sectioned 3 coverslips of slide 1 (AbdB AbdA Ubx).
  • sectioned 4 coverslips of centromere stained and counter-labeled nuclei.
This entry was posted in Summaries and tagged , , . Bookmark the permalink.